the equations of motion of projectile thrown in x-y plane from origin x=8t, y=6t-t then the angle of projection is -

Grade:11

2 Answers

Eshan
5 years ago
Dear student,

For projectile,x=u_xt=8t\implies u_x=8

=u_yt-\dfrac{1}{2}gt^2=6t-t^2
\implies u_y=6
Hence angle of projection=tan^{-1}(\dfrac{u_y}{u_x})=tan^{-1}(\dfrac{6}{8})=37^{\circ}
No name
one year ago
2uowgdukwsgkcuvssucgadsc6uktwdckygqedyucfeegg iuo1 f3yuerybufqberfuboyerui fyeulrbyfcjoweu yrxu2x3gxfouxybrexiorw3oujgxfujoegbxfjuowqgcfqu qegpcujfg wqeujcd gejoue gdqujee gcu2eg ofu12 y3cr7i13y31 7iry133ioiryowi73bryoixqebyor7uxqy